# Local expression of fractional corner charges in obstructed atomic insulators and its application to vertex-transitive polyhedra with arbitrary genus

### v1

@@ Abstract
In obstructed atomic insulators, fractional charges appear at the corners of the crystals in the shapes of vertex-transitive polyhedra, and are given by the filling anomaly divided by the number of corners. Recent studies reveal that the filling anomaly for the cases with genus 0 is universally given by the total charge at the Wyckoff position $1a$. In this study, we rewrite the formula in terms of the degree of sharpness of the corner, and show that the corner charge formula also holds for cases with arbitrary genus. We also extend our formula to vertex-transitive shell polyhedra, which are closed or open polyhedra without the bulk region, with all the vertices related by symmetry. Then, we show that the corner charges of such shell polyhedra are equal to the two-dimensional disclination charges of the corresponding disclinations. By identifying it with the disclination charge under the Wen-Zee action, we show that the coupling constant of the Wen-Zee action for a crystalline insulator is given by the total charge at the Wyckoff position at the disclination core.
@@ I Introduction
The study of topological phases of matter is one of the most important topics today [PhysRevB.76.045302, PhysRevB.27.6083, PhysRevLett.98.106803, PhysRevLett.95.146802, PhysRevLett.95.226801, PhysRevB.74.195312, doi:10.1126/science.1148047, doi:10.1126/science.1133734, PhysRevB.78.045426, PhysRevLett.106.106802, PhysRevB.96.245115, PhysRevB.91.161105, PhysRevB.95.081107, PhysRevB.90.165114, PhysRevX.7.041069, PhysRevB.98.081110, PhysRevLett.119.246402, Schindler_2018, doi:10.1126/sciadv.aat0346, PhysRevResearch.2.012067, PhysRevLett.124.036803, PhysRevResearch.2.043131, PhysRevX.11.041064, PhysRevB.105.045126, PhysRevB.109.085114, PhysRevB.111.155305, PhysRevB.99.245151, PhysRevB.103.205123, PhysRevResearch.1.033074, PhysRevB.102.165120, Po_2017, SI_SA_Watanabe, Bradlyn_2017, PhysRevB.97.035139, Aroyo:xo5013]. In the topologically nontrivial phases of matter, there are gapless excitations localized at their boundaries, which are robust against continuous deformations, and they are characterized by the classification indices called topological invariants. Since the proposal of Chern insulators [PhysRevB.76.045302, PhysRevB.27.6083] and $\mathbb{Z}_{2}$ topological insulators [PhysRevLett.98.106803, PhysRevLett.95.146802, PhysRevLett.95.226801, PhysRevB.74.195312, doi:10.1126/science.1148047, doi:10.1126/science.1133734], varieties of topological insulators have been discovered, such as topological crystalline insulators (TCIs) [PhysRevLett.106.106802, PhysRevB.96.245115, PhysRevB.91.161105, PhysRevB.95.081107, PhysRevB.90.165114, PhysRevX.7.041069] protected by crystallographic symmetries and higher-order topological insulators (HOTIs) [PhysRevB.98.081110, PhysRevLett.119.246402, Schindler_2018, doi:10.1126/sciadv.aat0346, PhysRevResearch.2.012067, PhysRevLett.124.036803, PhysRevResearch.2.043131] with gapless states on their hinges or corners. Furthermore, topological quantum chemistry (TQC) [Bradlyn_2017, PhysRevB.97.035139, Aroyo:xo5013, MTQC] and the theory of symmetry-based indicators [Po_2017, SI_SA_Watanabe] have allowed us to identify many topological materials [catalogue0, catalogue1, catalogue2, catalogue3, catalogue4, catalogue6].
@@
On the other hand, topologically trivial insulators, called atomic insulators (AIs), have been believed to have no topologically interesting features because their energy spectra are completely gapped including their boundaries. However, recent studies [PhysRevB.105.045126, PhysRevB.109.085114, PhysRevB.99.245151, PhysRevB.111.155305, PhysRevX.11.041064, PhysRevB.103.205123, PhysRevResearch.1.033074, PhysRevB.102.165120] have revealed that some AIs called obstructed atomic insulators (OAIs) have fractionally quantized charges on their corners, which are characterized by a topological invariant called a filling anomaly [PhysRevB.99.245151, PhysRevB.103.165109]. The filling anomaly is given by the total charge of a finite-sized system preserving crystallographic symmetries which connect all corners of the system, and the fractional corner charges can be obtained by equivalent distribution of the filling anomaly to the corners. The filling anomaly is well understood in two dimensions, and is given by the total charge at the Wyckoff position (WP) $1a$ located at the center of the systems [PhysRevB.103.205123, PhysRevB.102.165120].
@@
In previous works [PhysRevB.109.085114, PhysRevB.111.155305], we constructed a complete list of corner charge formulas for real and $k$ spaces in three-dimensional crystals. In this case, the crystal shapes need to be the vertex-transitive polyhedra with genus 0 [polyhedra, bridges2002:320, ROBERTSON197079, https://doi.org/10.1112/jlms/s2-2.1.125, IsogonalPrismatoids], and for all the corresponding crystal shapes for all space groups, we calculated the filling anomalies and corner charge formulas. As a result, the real-space formulas of the filling anomaly are universally given by the total charge at the bulk WP $1a$ located at the center of the crystal shapes, and the corner charge formulas are given by the filling anomaly divided by the number of corners, similar to the two-dimensional cases [PhysRevB.103.205123].
@@
Microscopically, we expect that the corner charge should be determined locally, i.e., it depends only on the shape of the corner. Thus we address the question of why the corner charge depends on the number of corners, which does not seem to be a local property. In this study, we derive a local expression of the corner charge formulas, and thereby, construct general corner charge formulas in crystals with the forms of the vertex-transitive polyhedra with arbitrary genus. Then we introduce the concept of vertex-transitive shell polyhedra (VTSPs), which are closed or open polyhedra without the bulk region, with all the vertices related by symmetry. Moreover, we construct the relationship between the two-dimensional disclination charges and the corner charges in the corresponding VTSPs, which leads to a general formulation of the disclination charges in two dimensions. Finally, we show that the coupling constant of the Wen-Zee action [PhysRevLett.69.953] in such crystalline insulators is given by the total charge at the Wyckoff position, where the disclination core is located.
@@ II Review of fractional corner charges in three-dimensional crystal shapes
In this section, we briefly review the fractional corner charges in three-dimensional crystal shapes based on the previous works [PhysRevB.109.085114, PhysRevB.111.155305], where we construct a complete list of corner charge formulas for all the SGs and all the corresponding crystal shapes with a quantized corner charge. Here, for clarity of explanation, we consider the corner charge formula for the SG No. 47 ($Pmmm$), which corresponds to the orthorhombic point-group symmetry $D_{2h}$. The unit cell and the corresponding crystal shape are shown in Fig. 1(a) and (b), respectively.
Figure 1: Calculation of the corner charge for the SG No. 47 (PmmmPmmm) in the previous study [PhysRevB.111.155305]. (a) Maximal Wyckoff positions in the SG No. 47. $\bm{a}_{i}\ (i=1,2,3)$ are primitive lattice vectors in the orthorhombic unit cell. The cyan cuboid represents the unit cell. (b) Crystal shape of a cuboid with orthorhombic symmetry consisting of $l\times m\times n$ unit cells. $x$, $y$, and $z$ axes are parallel to $\bm{a}_{1}$, $\bm{a}_{2}$, and $\bm{a}_{3}$, respectively.
@@ II.1 Preliminary
Next, we also review the filling anomaly [PhysRevB.99.245151, PhysRevB.103.165109]. In insulators of finite-sized crystals with some crystallographic symmetry, the charge neutrality for the whole crystal should be broken as long as the entire system is insulating, and the charge distribution fully reflects the crystallographic symmetry. In this case, the total charge of the entire system is called the filling anomaly. This filling anomaly is an integer and is a topological invariant. We note that the nontrivial filling anomaly is seen in certain AIs called OAIs, which are characterized by charge imbalance between electrons and ions at each WP, and we hereafter focus on OAIs.
@@
where $\alpha_{i}~(i=1,\cdots,7)$ are integer constants. Actually, $\alpha_{7}$ is proportional to the bulk charge density (per bulk unit cell), $\alpha_{i}~(i=4,5,6)$ are proportional to the surface charge densities (per surface unit cell) under the bulk charge neutrality condition $\alpha_{7}=0$. Under the bulk and surface charge neutrality conditions $\alpha_{i}=0~(i=4,\cdots,7)$, $\alpha_{i}~(i=1,2,3)$ are proportional to the hinge charge densities (per hinge unit cell). Hereafter, we assume that the bulk, surfaces, and hinges are all charge neutral. In this case, the filling anomaly $\eta_{l,m,n}$ is given by $\Delta a$ in terms of mod $2$. Since all corners of the cuboid are equivalent under the symmetry of the SG No. 47, the corner charge formula $Q_{\text{corner}}$ for the SG No. 47 is given by the filling anomaly divided by the number of corners [PhysRevB.111.155305],
@@ II.2 Real-space formulas of the corner charge in the vertex-transitive polyhedra with genus 0
In order to obtain the corner charge formulas for all the SGs, and all the corresponding crystal shapes, we should consider the crystal shapes in more detail. For the quantization of the corner charge, we need to restrict the crystal shapes to be the vertex-transitive polyhedra [polyhedra, bridges2002:320, ROBERTSON197079, https://doi.org/10.1112/jlms/s2-2.1.125, IsogonalPrismatoids], in which all of the corners are equivalent by crystallographic symmetries. In the previous works [PhysRevB.109.085114, PhysRevB.111.155305], we restricted ourselves to the vertex-transitive polyhedra with genus 0, and they are classified into the spherical family and the cylindrical family, whose examples are shown in Figs. 2(a) and (b), respectively. For both families, the corner charge formulas $Q_{\text{corner}}$ in the three-dimensional crystals are given by the filling anomaly $Q_{\text{tot}}$ divided by the number of vertices $N$ as follows:
@@
Here we used the result of our previous works [PhysRevB.109.085114, PhysRevB.111.155305] (Sec. II) that the filling anomaly $Q_{\text{tot}}$ is universally given by
@@
which is the total charge at the bulk WP $a$ located at the center of the crystal shapes. Here, the Wyckoff position $1a$ appears in the formula, because it is the WP located at the center of the crystal. We note that this result has the same form as in two dimensions [PhysRevB.103.205123], but unlike two dimensions, it is not known why the formula is same for all the SGs and for all the crystal shapes of vertex-transitive polyhedra.
@@ IV Corner charge formulas for arbitrary genus
Figure 4: Vertex-transitive polyhedron with genus $3$ under the tetrahedral group of rotations [Leopold].
@@ IV.1 Corner charge formula for three-dimensional polyhedra with arbitrary genus
The vertex-transitive polyhedra with higher genus are still largely unexplored mathematically, and the polyhedron shown in Fig. 4 is one of the examples introduced in Ref. [Leopold]. In such cases of higher genus, the following two points should be considered.
@@
Second, we revisit the definition of the filling anomaly. So far, we calculated the filling anomaly by counting the number of lattice points corresponding to the WP $a$ in the polyhedra. In this case, the Ehrhart polynominal [beck2015computing, DiazRobins1996] is a useful tool to count the lattice points in the lattice polyhedra.
Let $P$ be a lattice polyhedron, whose vertices are located at the lattice points. Let $tP$ be the $t$-fold dilation of $P$, where $t$ is any positive integer. In this case, the number of lattice points in the polyhedron $tP$ including the boundaries is given by the Ehrhart polynominal $L(t,P)$:
@@
where the coefficients $\alpha_{i}\ (i=0,1,2,3)$ are rational numbers. Here the coefficient $\alpha_{0}$ is given by
@@
where $\chi(P)$ is the Euler characteristic [beck2015computing, DiazRobins1996]. We note that since $P$ has the bulk region, the Euler characteristic $\chi(P)$ of $P$ is given by 1g1-g, not by $2-2g\ (=\chi(A))$.
@@
In particular, when the genus of $P$ is 0, Eq. (13) gives $\alpha_{0}=1$ and $|Q_{\text{total}}=-\Delta a\absolutevalue{e}$, as obtained in Eq. (5) [PhysRevB.109.085114, PhysRevB.111.155305]. This result is natural, by noting that the Ehrhart polynominal $L(t,P)$ can be defined even for $t=0$. In this case, for a polyhedron with genus 0, tPtP consists only of the origin, and $L(0,P)=\alpha_{0}=1$, which implies that corner charge formulas in three-dimensional crystal shapes with genus 0 are universally given by the total charge at their center.
@@ IV.2 Corner charge formula for two-dimensional shell polyhedra
Here the modulo part reflects the charge neutrality conditions for the surface charge density, because the VTSPs consist of the two-dimensional surfaces, and the charge neutrality conditions for hinges do not appear.
@@ IV.3 Example: Fractional corner charge in the two-dimensional shell polyhedron with genus 2
Figure 6: Calculation of the corner charge for the polyhedron $R$ in Fig. 5 with the SG No. 195. (a) Wyckoff positions in the SG No. 195. $\bm{a}_{i}\ (i=1,2,3)$ are primitive lattice vectors in the cubic unit cell. The cyan cube represents the unit cell. (b) Surface configuration of the polyhedron $R$ for the SG No. 195 when viewed from the [111] direction. Wyckoff position $1a$ is located at each corner of the triangular surface of $R$ (Fig. 5(c)).
@@
This result is natural from the following consideration. Because two vertices from two regular triangles (Fig. 6(b)) merge into one corner in the polyhedron $R$ in Fig. 5(c), we expect that the result of Eq. (20) equals a double of the corner charge formula of the two-dimensional crystal shapes with $C_{3}$ symmetry. It is indeed the case, since the corner charge of the two-dimensional crystal shapes with $C_{3}$ symmetry is given by $|-\frac{\Delta a}{3}\absolutevalue{e}$ in terms of modulo $|\absolutevalue{e}$ [PhysRevB.103.205123].
@@
Table 1: Summary for point groups of Wyckoff position $a$ and the filling anomalies for the 15 SGs. All the non-trivial filling anomalies are universally given by $2\Delta a$.
@@
Here we relate the corner charge of the crystal with the shape of VTSP in Sec. IVB to the two-dimensional fractional disclination charge discussed in Refs. [PhysRevB.99.245151, PhysRevB.101.115115]. We then show a general formulation of the disclination charges by using this relationship.

### v2

# Local expression of fractional corner charges in obstructed atomic insulators and relationship with the fractional disclination charges

@@ Abstract
In obstructed atomic insulators, fractionally quantized charges appear at the corners of the crystals in the shapes of vertex-transitive polyhedra, and are given by the filling anomaly divided by the number of corners. Recent studies reveal that the filling anomaly for the cases with genus 0 is universally given by the total charge at the Wyckoff position $1a$. In this study, we rewrite the formula in terms of the degree of sharpness of the corner, and show that the corner charge formula also holds for cases with arbitrary genus. We also extend our formula to vertex-transitive shell polyhedra, which are closed or open polyhedra without the bulk region, with all the vertices related by symmetry. Then, we show that the corner charges of such shell polyhedra are equal to the two-dimensional disclination charges of the corresponding disclinations. By identifying it with the disclination charge under the Wen-Zee action, we show that the coupling constant of the Wen-Zee action for a crystalline insulator is given by the total charge at the Wyckoff position at the disclination core.
@@ I Introduction
The study of topological phases of matter is one of the most important topics today [18, 44, 16, 23, 22, 17, 24, 6, 43, 19, 4, 14, 50, 37, 25, 45, 38, 36, 35, 1, 10, 21, 53, 29, 48, 49, 5, 39, 34, 51, 30, 52, 7, 9, 2, 46]. In the topologically nontrivial phases of matter, there are gapless excitations localized at their boundaries, which are robust against continuous deformations, and they are characterized by the classification indices called topological invariants. Since the proposal of Chern insulators [18, 44] and $\mathbb{Z}_{2}$ topological insulators [16, 23, 22, 17, 24, 6], varieties of topological insulators have been discovered, such as topological crystalline insulators (TCIs) [19, 4, 14, 50, 37, 25, 46] protected by crystallographic symmetries and higher-order topological insulators (HOTIs) [45, 38, 36, 35, 1, 10, 21] with gapless states on their hinges or corners. Furthermore, topological quantum chemistry (TQC) [7, 9, 2, 13] and the theory of symmetry-based indicators [30, 52] have allowed us to identify many topological materials [41, 56, 40, 47, 42, 55].
@@
On the other hand, topologically trivial insulators, called atomic insulators (AIs), have been believed to have no topologically interesting features because their energy spectra are completely gapped including their boundaries. However, recent studies [29, 48, 5, 49, 53, 39, 34, 51] have revealed that some AIs called obstructed atomic insulators (OAIs) have fractionally quantized charges on their corners, which are characterized by a topological invariant called a filling anomaly [5, 15]. The filling anomaly is given by the total charge of a finite-sized system preserving crystallographic symmetries which connect all corners of the system, and the fractional corner charges can be obtained by equivalent distribution of the filling anomaly to the corners. The filling anomaly is well understood in two dimensions, and is given by the total charge at the Wyckoff position (WP) $1a$ located at the center of the systems [39, 51].
@@
In previous works [48, 49], we constructed a complete list of corner charge formulas for real and $k$ spaces in three-dimensional crystals. In this case, the crystal shapes need to be the vertex-transitive polyhedra with genus 0 [11, 8, 33, 32, 20], and for all the corresponding crystal shapes for all space groups, we calculated the filling anomalies and corner charge formulas. As a result, the real-space formulas of the filling anomaly are universally given by the total charge at the bulk WP $1a$ located at the center of the crystal shapes, and the corner charge formulas are given by the filling anomaly divided by the number of corners, similar to the two-dimensional cases [39].
@@
Microscopically, we expect that the corner charge should be determined locally, i.e., it depends only on the shape of the corner. Thus we address the question of why the corner charge depends on the number of corners, which does not seem to be a local property. In this study, we derive a local expression of the corner charge formulas, and thereby, construct general corner charge formulas in crystals with the forms of the vertex-transitive polyhedra with arbitrary genus. Then we introduce the concept of vertex-transitive shell polyhedra (VTSPs), which are closed or open polyhedra without the bulk region, with all the vertices related by symmetry. Moreover, we construct the relationship between the two-dimensional disclination charges and the corner charges in the corresponding VTSPs, which leads to a general formulation of the disclination charges in two dimensions. Finally, we show that the coupling constant of the Wen-Zee action [54] in such crystalline insulators is given by the total charge at the WP, where the disclination core is located. We note that near the surfaces of the crystals, the electronic states and nuclear positions may shift slightly from the ideal configuration of crystals. Nevertheless, the fractional corner charge survives such deviation from perfect periodicity, as long as such deviation does not break the focused point-group symmetries, and the lattice periodicity along the hinges and the surfaces. Thus, the fractional corner charge is robust against those surface effects, where the lattice periodicity in the bulk is preserved along the hinges and the surfaces.
@@ II Review of fractional corner charges in three-dimensional crystal shapes
In this section, we briefly review the fractional corner charges in three-dimensional crystal shapes based on the previous works [48, 49], where we construct a complete list of corner charge formulas for all the SGs and all the corresponding crystal shapes with a quantized corner charge. Here, for clarity of explanation, we consider the corner charge formula for the SG No. 47 ($Pmmm$), which corresponds to the orthorhombic point-group symmetry $D_{2h}$. The unit cell and the corresponding crystal shape are shown in Fig. 1(a) and (b), respectively.
Figure 1: Calculation of the corner charge for the SG No. 47 (PmmmPmmm) in the previous study [49]. (a) Maximal WPs in the SG No. 47. $\bm{a}_{i}\ (i=1,2,3)$ are primitive lattice vectors in the orthorhombic unit cell. The cyan cuboid represents the unit cell. (b) Crystal shape of a cuboid with orthorhombic symmetry consisting of $l\times m\times n$ unit cells. $x$, $y$, and $z$ axes are parallel to $\bm{a}_{1}$, $\bm{a}_{2}$, and $\bm{a}_{3}$, respectively.
@@ II.1 Preliminary
Next, we also review the filling anomaly [5, 15]. In insulators of finite-sized crystals with some crystallographic symmetry, the charge neutrality for the whole crystal should be broken as long as the entire system is insulating, and the charge distribution fully reflects the crystallographic symmetry. In this case, the total charge of the entire system is called the filling anomaly. This filling anomaly is an integer and is a topological invariant. We note that the nontrivial filling anomaly is seen in certain AIs called OAIs, which are characterized by charge imbalance between electrons and ions at each WP, and we hereafter focus on OAIs.
@@
where $\alpha_{i}~(i=1,\cdots,7)$ are integer constants. Actually, $\alpha_{7}$ is proportional to the bulk charge density (per bulk unit cell), $\alpha_{i}~(i=4,5,6)$ are proportional to the surface charge densities (per surface unit cell) under the bulk charge neutrality condition $\alpha_{7}=0$. Under the bulk and surface charge neutrality conditions $\alpha_{i}=0~(i=4,\cdots,7)$, $\alpha_{i}~(i=1,2,3)$ are proportional to the hinge charge densities (per hinge unit cell). Hereafter, we assume that the bulk, surfaces, and hinges are all charge neutral. In this case, the filling anomaly $\eta_{l,m,n}$ is given by $\Delta a$ in terms of mod $2$. Since all corners of the cuboid are equivalent under the point-group symmetry $D_{2h}$, which is the local symmetry at the WP $a$ in the SG No. 47, the corner charge formula $Q_{\text{corner}}$ for the SG No. 47 is given by the filling anomaly divided by the number of corners [49],
@@
where we explicitly write the electron charge $|-\absolutevalue{e}$.
@@ II.2 Real-space formulas of the corner charge in the vertex-transitive polyhedra with genus 0
In order to obtain the corner charge formulas for all the SGs, and all the corresponding crystal shapes, we should consider the crystal shapes in more detail. For the quantization of the corner charge, we need to restrict the crystal shapes to be the vertex-transitive polyhedra [11, 8, 33, 32, 20], in which all of the corners are equivalent by crystallographic symmetries, and thus the filling anomaly $Q_{\text{tot}}$ is divided equally among their corners. In the previous works [48, 49], we restricted ourselves to the vertex-transitive polyhedra with genus 0, and they are classified into the spherical family and the cylindrical family, whose examples are shown in Figs. 2(a) and (b), respectively. For both families, the corner charge formulas $Q_{\text{corner}}$ in the three-dimensional crystals are given by the filling anomaly $Q_{\text{tot}}$ divided by the number of vertices $N$ as follows:
@@
Here we used the result of our previous works [48, 49] (Sec. II) that the filling anomaly $Q_{\text{tot}}$ is universally given by
@@
which is the total charge at the bulk WP $a$ located at the center of the crystal shapes. Here, the WP $1a$ appears in the formula, because it is the WP located at the center of the crystal. We note that this result has the same form as in two dimensions [39], but unlike two dimensions, it is not known why the formula is same for all the SGs and for all the crystal shapes of vertex-transitive polyhedra.
@@ IV Corner charge formulas for arbitrary genus
Figure 4: Vertex-transitive polyhedron with genus $3$ under the tetrahedral group of rotations [26].
@@ IV.1 Corner charge formula for three-dimensional polyhedra with arbitrary genus
The vertex-transitive polyhedra with higher genus are still largely unexplored mathematically, and the polyhedron shown in Fig. 4 is one of the examples introduced in Ref. [26]. In such cases of higher genus, the following two points should be considered.
@@
Second, we revisit the definition of the filling anomaly. So far, we calculated the filling anomaly by counting the number of lattice points corresponding to the WP $a$ in the polyhedra. In this case, the Ehrhart polynominal [3, 12] is a useful tool to count the lattice points in the lattice polyhedra.
Let $P$ be a lattice polyhedron, in which all the vertices are located at the integer lattice points. Let $tP$ be the $t$-fold dilation of $P$, where $t$ is any positive integer. In this case, the number of lattice points in the polyhedron $tP$ including the boundaries is given by the Ehrhart polynominal $L(t,P)$:
@@
where the coefficients $\alpha_{i}\ (i=0,1,2,3)$ are rational numbers. Here the coefficient $\alpha_{0}$ is given by [3, 12]
@@
where $\chi(P)$ is the Euler characteristic. We note that since $P$ has the bulk region, the Euler characteristic $\chi(P)$ of $P$ is given by 1g1-g, not by $2-2g\ (=\chi(A))$.
@@
In particular, when the genus of $P$ is 0, Eq. (13) gives $\alpha_{0}=1$ and $|Q_{\text{total}}=-\Delta a\absolutevalue{e}$, as obtained in Eq. (5) [48, 49]. This result is natural, by noting that the Ehrhart polynominal $L(t,P)$ can be defined even for $t=0$. In this case, for a polyhedron with genus 0, tPtP consists only of the origin, and $L(0,P)=\alpha_{0}=1$, which implies that corner charge formulas in three-dimensional crystal shapes with genus 0 are universally given by the total charge at their center.
@@ IV.2 Corner charge formula for two-dimensional shell polyhedra
The ambiguity of the corner charge in terms of modulo $|\absolutevalue{e}$ in Eq. (16) comes from ”decoration” near the corners, i.e., deviation of the nuclei positions and the electronic states near the corners away from those in the bulk. We note that there might also be decorations near the hinges, which can modify the modulo part, but here we assume them for a while. This assumption is convenient in the comparison between the corner charge and the disclination charge in the next section. We will discuss the effect of the decorations near the hinges in Conclusion.
@@ IV.3 Example: Fractional corner charge in the two-dimensional shell polyhedron with genus 2
Figure 6: Calculation of the corner charge for the polyhedron $R$ in Fig. 5 with the SG No. 195. (a) WPs in the SG No. 195. $\bm{a}_{i}\ (i=1,2,3)$ are primitive lattice vectors in the cubic unit cell. The cyan cube represents the unit cell. (b) Surface configuration of the polyhedron $R$ for the SG No. 195 when viewed from the [111] direction. WP $1a$ is located at each corner of the triangular surface of $R$ (Fig. 5(c)).
@@
This result is natural from the following consideration. Because two vertices from two regular triangles (Fig. 6(b)) merge into one corner in the polyhedron $R$ in Fig. 5(c), we expect that the result of Eq. (20) equals a double of the corner charge formula of the two-dimensional crystal shapes with $C_{3}$ symmetry. It is indeed the case, since the corner charge of the two-dimensional crystal shapes with $C_{3}$ symmetry is given by $|-\frac{\Delta a}{3}\absolutevalue{e}$ in terms of modulo $|\absolutevalue{e}$ [39].
@@
Table 1: Summary for point groups of WP $a$ and the filling anomalies for the 15 SGs. All the non-trivial filling anomalies are universally given by $2\Delta a$.
@@
Here we relate the corner charge of the crystal with the shape of VTSP in Sec. IVB to the two-dimensional fractional disclination charge discussed in Refs. [5, 27]. We then show a general formulation of the disclination charges by using this relationship.